# Sim-to-Real Transfer of Deep Reinforcement Learning Agents for Online Coverage Path Planning

### v2

@@ Abstract
Sim-to-real transfer presents a difficult challenge, where models trained in simulation are to be deployed in the real world. The distribution shift between the two settings leads to biased representations of the dynamics, and thus to suboptimal predictions in the real-world environment. In this work, we tackle the challenge of sim-to-real transfer of reinforcement learning (RL) agents for coverage path planning (CPP). In CPP, the task is for a robot to find a path that covers every point of a confined area. Specifically, we consider the case where the environment is unknown, and the agent needs to plan the path online while mapping the environment. We bridge the sim-to-real gap through a semi-virtual environment, including a real robot and real-time aspects, while utilizing a simulated sensor and obstacles to enable environment randomization and automated episode resetting. We investigate what level of fine-tuning is needed for adapting to a realistic setting, comparing to an agent trained solely in simulation. We find that a high inference frequency allows first-order Markovian policies to transfer directly from simulation, while higher-order policies can be fine-tuned to further reduce the sim-to-real gap. Moreover, they can operate at a lower frequency, thus reducing computational requirements. In both cases, our approaches transfer state-of-the-art results from simulation to the real domain, where direct learning would take in the order of weeks with manual interaction, that is, it would be completely infeasible.
Introduction
A practical limitation for training machine learning models is the need for large amounts of real-world data, which is tedious and time-consuming to collect. In particular, training reinforcement learning (RL) agents for robotics from scratch requires access to the robot for the full duration of the training process. Meanwhile, during the early training phase, the agent is more likely to make mistakes that may damage the hardware or require human supervision for resetting the episodes. Instead, learning in simulation presents an attractive alternative (Jonnarth, Zhao, and Felsberg 2024). However, due to differences in the dynamics between the simulation and the real world, transferring a model from simulation to reality is challenging. Prior work reduce the sim-to-real gap by improving the simulation, e.g. by injecting realistic noise (Kaufmann et al. 2023; Zhao et al. 2020), applying domain randomization (Muratore et al. 2021; Tobin et al. 2017), or through meta learning (Arndt et al. 2020; Nagabandi et al. 2019). Our goal is to transfer, for the task of coverage path planning (CPP), the state-of-the-art performance of models trained in simulation to real environments, by fine-tuning in a realistic setting.
@@ Introduction
In CPP, the task is to find a path that covers all of the free space of an environment. If the environment is known, an optimal path can be planned offline (Huang 2001). If it is unknown, the path has to be planned online during mapping of the environment, e.g. by a robot, and an optimal path cannot be found in the general case (Galceran and Carreras 2013). CPP has found its uses in many robotic applications, such as lawn mowing (Cao, Huang, and Hall 1988), vacuum cleaning (Yasutomi, Yamada, and Tsukamoto 1988), search-and-rescue (Jia et al. 2016), and exploration (Xu et al. 2022).
@@
When training reinforcement learning models in real time on physical robots, additional considerations need to be accounted for compared to training in simulation. (1) There exists a mismatch between the real and simulated robot kinematics, dynamics, and sensing, such as slippage and noisy localization. This leads to different transition probabilities in the real world, where the predicted actions based on simulation may be suboptimal. (2) Due to inertia and latencies in the system, the dynamics are non-Markovian (Haarnoja et al. 2018b). This violates the common assumption that the Markov property holds. (3) Since the robot keeps moving during the various computations in the training process, the state observed by the agent is not perfectly aligned with the true state. This introduces a delay, where the agent predicts actions based on outdated observations of the environment. In this work, we address these problems.
@@
To smoothen the transition into the real world, we use a real robot in a semi-virtual environment, utilizing a highly accurate positioning system in an indoor planar setting, with a simulated sensor and obstacles. This introduces the previously mentioned real-time aspects of reinforcement learning in a controlled manner, while allowing flexibility in utilizing automated episode resetting and varying training environments without having to physically construct them. By fine-tuning and evaluating CPP policies in this semi-virtual setting, we can draw conclusions about how we expect similar policies to generalize to fully realistic environments.
@@
To reduce latency in the training process, we perform model updates in parallel with the action and state selection (Yuan and Mahmood 2022), and perform all computations on on-board hardware. We utilize soft actor-critic learning (Haarnoja et al. 2018a) for its sample efficiency and its efficacy on continuous action spaces. To account for the non-Markovian dynamics in a lightweight manner, we include past actions in the observation space. Finally, to reduce the mismatch between the simulated and real kinematics, we measure the real-world linear and angular accelerations as well as action delay, and incorporate them into the simulation. We find that a high inference frequency enables first-order Markovian policies to transfer to a real setting. By introducing delays and action observations, higher order Markovian models can be fine-tuned to further reduce the sim-to-real gap. Moreover, these models can operate at a lower frequency, thus reducing the computational requirements for deployed systems.
Our contributions can be summarized as follows:
  • We propose to divide the sim-to-real problem into two steps with an intermediate case of a real robot in a virtual environment.
  • We show that this division enables a transfer of state-of-the-art RL policies for CPP from simulation to the real robot.
@@
  • We propose to perform data collection and model updates in parallel, which enables real-time fine-tuning online without a separate system or stopping the robot for model updates.
  • We evaluate the effects of inference frequency and fine-tuning on the performance of the real robot.
Related Work
@@
This paper relates to coverage path planning, transferring models from simulation to the real world, and online RL in real time. We summarize the related work below.
@@
Coverage path planning methods can roughly be categorized as planning-based or learning-based. Planning-based methods include decomposition methods, which divide the area into cells based on e.g. boustrophedon cellular decomposition (BCD) (Choset and Pignon 1998) or Morse decomposition (Acar et al. 2002), where each cell is covered with a pre-defined pattern. Grid-based methods, such as Spiral-STC (Gabriely and Rimon 2002) and the backtracking spiral algorithm (BSA) (Gonzalez et al. 2005), discretize the area into even smaller cells, e.g. squares, with a similar size to the agent. Then, a path is planned based on motion primitives connecting adjacent cells, i.e. to move up, down, left, or right on the grid, such that each cell is visited at least once. Frontier-based methods plan a path to a chosen point on the frontier, i.e. on the boundary between covered and non-covered regions. The choice of frontier point can be based on different aspects, such as the distance to the agent (Yamauchi 1997), the path of a rapidly exploring random tree (RRT) (Umari and Mukhopadhyay 2017) or the gradient in a potential field (Yu et al. 2021). Learning-based methods apply machine learning techniques, typically in combination with planning-based methods, to find coverage paths. Reinforcement learning is the most popular paradigm due to the sequential nature of the task. Chen et al. (2019) use RL to find the order in which to cover the cells generated by BCD. Discrete methods (Piardi et al. 2019; Kyaw et al. 2020) use RL to learn which motion primitives to perform. RL has also been combined with frontier-based methods, either for predicting the cost of each frontier point (Niroui et al. 2019), or for learning the control signals to navigate to a chosen point (Hu et al. 2020). In contrast, Jonnarth, Zhao, and Felsberg (2024) propose to learn continuous control signals end-to-end from a built map and sensor data. In our work, we adopt this approach to transfer agents from simulation to a real setting. To the best of our knowledge, we are the first to do so for end-to-end RL policies in CPP.
@@
Sim-to-real transfer. Transferring from simulation to the real world is challenging due to mismatch in both sensing and actuation (Zhao, Queralta, and Westerlund 2020). Prior work has approached this challenge by different means. Domain randomization has been utilized to randomize physical parameters in simulation, such as mass and joint damping (Muratore et al. 2021), or to randomize textures and lighting in the image domain (Tobin et al. 2017). Other works introduce perturbations in the sensing and actuation (Zhao et al. 2020). Meta learning methods aim to quickly adapt to new unseen tasks from a wide variety of training task, such as adapting to the real world from simulation (Arndt et al. 2020; Nagabandi et al. 2019). Another approach is to learn from expert demonstrations through imitation learning (Yan et al. 2017), which has previously been applied to coverage path planning (Hu et al. 2020). When it comes to robot control, Niroui et al. (2019) deploy an RL policy for CPP trained in simulation on a differential drive robot without fine-tuning. The policy predicts the next frontier node, where a separate non-learned module navigates to it, thus being less affected by misaligned kinematics between simulation and reality. Kaufmann et al. (2023) transfer a lightweight RL control policy using a pre-trained perception system for first-person-view drone racing, utilizing empirical noise models to improve the fidelity of the simulator. They collect perception and dynamics residuals in the real world based on a highly accurate positioning system, which they use to augment the simulation. In contrast to these works, we fine-tune our model online in the real world.
@@
Online RL. Compared to turn-based and simulated environments, where the time for action selection and policy updates are assumed to be zero and the next state can be advanced to in an instant, online RL in the real world presents new challenges where these assumptions do not hold. The core idea in the literature for accounting for this discrepancy is to parallelize aspects of the environment interactions and the model training. Bakker et al. (2006) explore quasi-online reinforcement learning, where a model of the environment is built online and in parallel to training the policy based on the built environment model. Ramstedt and Pal (2019) propose real-time reinforcement learning, where the action selection occurs in parallel with the state selection. Concretely, given the current state and action to be taken, the agent and environment compute the next set of state and action concurrently. This approach takes into account the time for action selection by delaying the observed state by one time step. However, it does not consider the time for model updates, which is typically much larger. Other works (Haarnoja et al. 2019; Wang, Vasan, and Mahmood 2023) distribute the model training for soft actor-critic (SAC) (Haarnoja et al. 2018a) to a separate system, while performing only the lightweight action and state selections on the on-board target hardware. This allows the data collection to be executed independently from the model updates, where the policy and replay buffer are periodically synchronized. Yuan and Mahmood (2022) employ a similar approach, but they run everything on a single edge device. They divide the training process into separate processes for the action selection, state selection, batch sampling, and gradient computation. We follow this approach. However, in our experimental setup, the action selection and batch sampling were much faster than the state selection and gradient computations, so we only use two threads in order to reduce complexity and communication overhead.
Background
In this section, we first formulate the CPP problem as a partially observable Markov decision process, and then briefly describe the approach by Jonnarth, Zhao, and Felsberg (2024), which we use to train RL agents for CPP.
Problem Formulation
We aim to transfer an RL agent trained in simulation to the real world, in the task of online coverage path planning in unknown environments. The task is for the agent to simultaneously map the a priori unknown environment geometry, while finding a path that covers all of its free space. A point in the free space is considered covered when the distance to the center of the agent is less than the coverage radius $d$, and the point is in the field-of-view of the agent. The coverage radius may be smaller or equal to the agent radius $r$ as in the lawn mowing problem, or larger than $r$ as in the exploration problem. For mapping the environment, the agent is equipped with a ranging sensor, which is used to add obstacle detections to a global map.
@@
We formulate the task as a partially observable Markov decision process (POMDP), consisting of an agent that predicts actions in an environment such as to maximize a reward function. In discrete time steps $t\in\{1,...,T\}$, the agent predicts actions $a_{t}\sim\pi(a_{t}|o_{t})$ according to policy $\pi$, based on observations $o_{t}$ of the state $s_{t}$. The policy is a neural network that predicts continuous control signals $a_{t}$ for the agent. Subsequently, the agent receives a new observation and reward $o_{t},r_{t}\sim p(o_{t},r_{t}|s_{t},a_{t-1})$ from the environment, where $s_{t}\sim p(s_{t}|s_{t-1},a_{t-1})$. The goal for the agent is to maximize the expected discounted reward $\mathbb{E}(\sum_{t=1}^{T}\gamma^{t}r_{t})$, with discount factor $\gamma$.
@@
In our particular problem, the state $s_{t}$ includes the full environment geometry with all obstacles, the set of points that have been covered, as well as the pose of the agent. However, the agent only has information about what it has observed, so the observation $o_{t}$ consists of the part of the environment that has been mapped until time step $t$, along with covered points, agent pose, and sensor data. In the setting with a real robot, the first order Markovian property is not fulfilled, due to the dynamics including inertia and momentum. The usual approach to augment previous states in the current one becomes infeasible if the motion is implicitly represented by the agent-centered environment maps. Thus higher-order effects lead to model errors that need correction in subsequent steps.
Reinforcement Learning of Coverage Paths
In this section, we summarize the approach presented by Jonnarth, Zhao, and Felsberg (2024) for learning CPP with RL, in terms of observation space, action space, reward function, and model architecture.
Map observation. In order to represent the environment, covered regions and detected obstacles are recorded in the form of global grid-like 2D maps. To represent large regions as input to the agent, multiple local maps are used with varying scales, each with the same grid size. The maps contain increasingly larger crops of the local area, with decreasing resolution. The finest scale shows detailed information near the agent, while the coarsest scale contains the largest area at a low resolution. In addition to coverage and obstacle maps, multi-scale frontier maps are also recorded. The frontier is the boundary between covered and non-covered space, which allows the agent to easily plan paths towards non-covered regions. All maps are represented as egocentric maps, that is, they are aligned such that the agent is in the center facing upwards.
Sensor observation. To detect obstacles, the agent is equipped with a simulated light detection and ranging (lidar) sensor, which measures distances in fixed angles relative to the agent. When obstacles are detected, they are added to the obstacle map based on the known global pose of the agent. The lidar distances are included in the observation space after normalization to $[0,1]$.

### v3

@@ Abstract
Coverage path planning (CPP) is the problem of finding a path that covers the entire free space of a confined area, with applications ranging from robotic lawn mowing to search-and-rescue. While for known environments, offline methods can find provably complete paths, and in some cases optimal solutions, unknown environments need to be planned online during mapping. We investigate the suitability of continuous-space reinforcement learning (RL) for this challenging problem, and propose a computationally feasible egocentric map representation based on frontiers, as well as a novel reward term based on total variation to promote complete coverage. Compared to existing classical methods, this approach allows for a flexible path space, and enables the agent to adapt to specific environment characteristics. Meanwhile, the deployment of RL models on real robot systems is difficult. Training from scratch may be infeasible due to slow convergence times, while transferring from simulation to reality, i.e. sim-to-real transfer, is a key challenge in itself. We bridge the sim-to-real gap through a semi-virtual environment, including a real robot and real-time aspects, while utilizing a simulated sensor and obstacles to enable environment randomization and automated episode resetting. We investigate what level of fine-tuning is needed for adapting to a realistic setting. Through extensive experiments, we show that our approach surpasses the performance of both previous RL-based approaches and highly specialized methods across multiple CPP variations in simulation. Meanwhile, our method successfully transfers to a real robot. Our code implementation can be found online. [^1] Link to code repository: https://github.com/arvijj/rl-cpp
Index Terms:
Coverage path planning, end-to-end learning, exploration, online, real-time, reinforcement learning, robotics, sim-to-real transfer, total variation.
[…]
=-21pt
Refer to caption Refer to caption
Figure 1: Learned paths for exploration (left) and lawn mowing (right), including the start (red triangle) and end position (green square).
I Introduction
@@
Task automation is an ever-growing field, and as automated systems become more intelligent, they are expected to solve increasingly challenging problems. One such problem is coverage path planning (CPP), where the task is to find a path that covers all of the free space of an environment. If the environment is known, an optimal path can be planned offline [1]. In contrast, if it is unknown, the path has to be planned online during mapping of the environment, e.g. by a robot, and an optimal path cannot be found in the general case [2]. CPP has found its uses in many robotic applications, such as lawn mowing [3], vacuum cleaning [4], search-and-rescue [5], and exploration [6].
We aim to develop a task- and robot-agnostic method for online CPP in unknown environments. This is a challenging task, as intricacies in certain applications may be difficult to model, or even impossible, in some cases. In the real world, unforeseen discrepancies between the modeled and true environment may occur, e.g. due to wheel slip. Incompleteness of the present model must be assumed to account for open world scenarios. Thus, we approach the problem from the perspective of learning through embodiment. Reinforcement learning (RL) lends itself as a natural choice, where an agent can interact with the world and adapt to specific environment conditions, without the need for an explicit model. Concretely, we consider the case where an RL model directly predicts continuous control signals for an agent from sensor data. We present an RL approach for the CPP problem in terms of a multi-scale map feature representation using frontiers, a continuous action space, a network architecture based on grouped convolutions, and a reward term based on total variation to encourage complete coverage. Our approach finds coverage paths in complex environments for different variations of the CPP problem, as can be seen in Fig. 1.
A practical limitation for training machine learning models is the need for large amounts of real-world data, which is tedious and time-consuming to collect. In particular, training reinforcement learning agents for robotics from scratch requires access to the robot for the full duration of the training process. Meanwhile, during the early training phase, the agent is more likely to make mistakes that may damage the hardware or require human supervision for resetting the episodes. Instead, learning in simulation presents an attractive alternative. However, due to differences in the dynamics between the simulation and the real world, transferring a model from simulation to reality is challenging. Prior work reduce the sim-to-real gap by improving the simulation, e.g. by injecting realistic noise [7, 8], applying domain randomization [9, 10], or through meta learning [11, 12]. Our goal is to transfer, for the task of coverage path planning, models trained in simulation to real environments, by fine-tuning in a realistic setting.
@@
When training reinforcement learning models in real time on physical robots, additional considerations need to be accounted for compared to training in simulation. (1) There exists a mismatch between the real and simulated robot kinematics, dynamics, and sensing, such as slippage and noisy localization. This leads to different transition probabilities in the real world, where the predicted actions based on simulation may be suboptimal. (2) Due to inertia and latencies in the system, the dynamics are non-Markovian [13]. This violates the common assumption that the Markov property holds. (3) Since the robot keeps moving during the various computations in the training process, the state observed by the agent is not perfectly aligned with the true state. This introduces a delay, where the agent predicts actions based on outdated observations of the environment. In this work, we address these problems.
@@
To smoothen the transition into the real world, we use a real robot in a semi-virtual environment, utilizing a highly accurate positioning system in an indoor planar setting, with a simulated sensor and obstacles. This introduces the previously mentioned real-time aspects of reinforcement learning in a controlled manner, while allowing flexibility in utilizing automated episode resetting and varying training environments without having to physically construct them. The conclusions drawn from the fine-tuning experiments in this semi-virtual setting are also relevant for the case of completely real environments, although they might need to be complemented with perception-specific findings.
@@
To reduce latency in the training process, we perform model updates in parallel with the action and state selection [14], and perform all computations on on-board hardware. We utilize soft actor-critic learning [15] for its sample efficiency and its efficacy on continuous action spaces. To account for the non-Markovian dynamics in a lightweight manner, we include past actions in the observation space. To reduce the mismatch between the simulated and real kinematics, we measure the real-world linear and angular accelerations as well as action delay, and incorporate them into the simulation. We find that a high inference frequency enables first-order Markovian policies to transfer to a real setting. By introducing delays and action observations, higher-order Markovian models can be fine-tuned to further reduce the sim-to-real gap. Moreover, these models can operate at a lower frequency, thus reducing computational requirements for deployed systems.
This work builds on the conference paper [16], and is part of the PhD thesis [17]. Our contributions are as follows:
  • We propose an end-to-end deep reinforcement learning approach in continuous space for the online CPP problem. This includes a multi-scale map representation exploiting frontiers, a network architecture using grouped convolutions, and a novel reward term based on total variation.
@@
  • We propose to divide the sim-to-real problem into two steps with an intermediate case of a real robot in a virtual environment. By performing data collection and model updates in parallel, our approach enables sim-to-real transfer of CPP policies through real-time fine-tuning online without a separate system or stopping the robot for model updates.
  • Finally, we conduct extensive experiments, both in simulation and in a realistic setting, including detailed ablations on our proposed map representation, network architecture, and reward function, all of which improve the coverage time. Our approach surpasses the performance of both classical and RL-based methods, while successfully transferring to a real robot.
II Related Work
@@
This work relates to coverage path planning, transferring models from simulation to the real world, and online RL in real time. We summarize the related work below.
II-A Coverage Path Planning
@@
Coverage path planning methods can roughly be categorized as planning-based or learning-based. Planning-based methods include decomposition methods, which divide the area into cells based on e.g. boustrophedon cellular decomposition (BCD) [18] or Morse decomposition [19], where each cell is covered with a pre-defined pattern. Grid-based methods, such as Spiral-STC [20] and the backtracking spiral algorithm (BSA) [21], discretize the area into even smaller cells, e.g. squares, with a similar size to the agent. Then, a path is planned based on motion primitives connecting adjacent cells, i.e. to move up, down, left, or right on the grid, such that each cell is visited at least once. Frontier-based methods plan a path to a chosen point on the frontier, i.e. on the boundary between covered and non-covered regions. The choice of frontier point can be based on different aspects, such as the distance to the agent [22], the path of a rapidly exploring random tree (RRT) [23] or the gradient in a potential field [24]. Learning-based methods apply machine learning techniques, typically in combination with planning-based methods, to find coverage paths. Reinforcement learning is the most popular paradigm due to the sequential nature of the task. Chen et al. [25] use RL to find the order in which to cover the cells generated by BCD. Discrete methods [26, 27] use RL to learn which motion primitives to perform. RL has also been combined with frontier-based methods, either for predicting the cost of each frontier point [28], or for learning the control signals to navigate to a chosen point [29]. In contrast, we learn continuous control signals end-to-end from a built map and sensor data to fully utilize the flexibility of RL. To the best of our knowledge, we are the first to do so for CPP.
II-B Map Representation
To perform coverage path planning, the environment needs to be represented in a suitable manner. Similar to previous work, we discretize the map into a 2D grid with sufficiently high resolution to accurately represent the environment. This map-based approach presents different choices for the input feature representation. Saha et al. [30] observe such maps in full resolution, where the effort for a $d\times d$ grid is $\mathcal{O}(d^{2})$, which is infeasible for large environments. Niroui et al. [28] resize the maps to keep the input size fixed. However, this means that the information in each grid cell varies between environments, and hinders learning and generalization for differently sized environments. Meanwhile, Shrestha et al. [31] propose to learn the map in unexplored regions. Instead of considering the whole environment at once, other works [32, 33] observe a local neighborhood around the agent. While the computational cost is manageable, the long-term planning potential is limited as the space beyond the local neighborhood cannot be observed. For example, if the local neighborhood is fully covered, the agent must pick a direction at random to explore further. To avoid the aforementioned limitations, we use multiple maps in different scales, similar to Klamt and Behnke [34].
II-C Sim-to-real Transfer
@@
Transferring from simulation to the real world is challenging due to mismatch in both sensing and actuation [35]. Prior work has approached this challenge by different means. Domain randomization has been utilized to randomize physical parameters in simulation, such as mass and joint damping [9], or to randomize textures and lighting conditions in the image domain [10]. Other works introduce perturbations in the sensing and actuation [8]. The motivation behind these approaches is that a highly randomized simulation would cover the real-world distribution, while avoiding a highly accurate model of it [35]. Instead of randomizing various aspects in simulation, we train directly on the real distribution by fine-tuning in the real world. Meta learning methods aim to quickly adapt to new unseen tasks from a wide variety of training task, such as adapting to the real world from simulation [11, 12]. Another approach is to learn from expert demonstrations through imitation learning [36], which has previously been applied to coverage path planning [29]. In contrast, we avoid human demonstrations in order to reduce the manual effort. When it comes to robot control, Niroui et al. [28] deploy an RL policy for CPP trained in simulation on a differential drive robot without fine-tuning. The policy predicts the next frontier node, where a separate non-learned module navigates to it, thus being less affected by misaligned kinematics between simulation and reality. Kaufmann et al. [7] transfer a lightweight RL control policy using a pre-trained perception system for first-person-view drone racing, utilizing empirical noise models to improve the fidelity of the simulator. They collect perception and dynamics residuals in the real world based on a highly accurate positioning system, which they use to augment the simulation. In contrast to these works, we fine-tune our model online in the real world.
II-D Real-time Reinforcement Learning
@@
Compared to turn-based and simulated environments, where the time for action selection and policy updates are assumed to be zero and the next state can be advanced to instantly, RL in the real world presents new challenges where these assumptions do not hold. For example, the forward and backward passes of the policy network during action selection and model updates take time, which cannot be neglected. To take this into account, the core idea in the literature is to parallelize aspects of the environment interactions and the model training. Bakker et al. [37] explore quasi-online reinforcement learning, where a model of the environment is built online and in parallel to training the policy based on the built environment model. Ramstedt and Pal [38] propose to perform the action selection in parallel with the state selection. Concretely, given the current state and action to be taken, the next state and new action are computed concurrently. This approach takes into account the time for action selection by delaying the observed state by one time step. However, it does not consider the time for model updates, which is typically much larger. Other works [39, 40] distribute the model training for soft actor-critic (SAC) [15] to a separate system, while performing only the lightweight action and state selections on the on-board target hardware. This allows the data collection to be executed independently from the model updates, where the policy and replay buffer are periodically synchronized. Within this paradigm, Wang et al. [40] propose ReLoD, a system to efficiently distribute the learning algorithm over a resource-limited local computer and a powerful remote computer. Yuan and Mahmood [14] employ a similar approach, but they run everything on a single edge device. They divide the training process into separate processes for the action selection, state selection, batch sampling, and gradient computation. We follow this approach. However, in our experimental setup, the action selection and batch sampling were much faster than the state selection and gradient computations, so we only use two threads in order to reduce complexity and communication overhead.
III Reinforcement Learning of Coverage Paths
This section presents our approach for learning coverage paths using reinforcement learning. First, we define the online CPP problem in Section III-A, and subsequently formulate it as a partially observable Markow decision process (POMDP) in Section III-B. After that, we present our RL-based approach in terms of observation space in Section III-C, action space in Section III-D, reward function in Section III-E, and agent architecture in Section III-F.
III-A Problem Definition and Delineations
The goal is to navigate with an agent of radius $r$ to visit all accessible points in a confined area without prior knowledge of its geometry. The free space includes all points inside the area that are not occupied by obstacles. A point is considered visited when the distance between the point and the agent is less than the coverage radius $d$, and the point is within the field-of-view of the agent. This definition unifies variations where $d\leq r$, which we define as the lawn mowing problem, and where $d>r$, which we refer to as exploration. To interactively gain knowledge about the environment, and for mapping its geometry, the agent needs some form of sensor. Both ranging sensors [29] and depth cameras [41] have been utilized for this purpose. Following Xu et al. [6], we choose a simulated 2D light detection and ranging (lidar) sensor that can detect obstacles in fixed angles relative to the agent, although our proposed framework is not limited to this choice. Based on the pose of the agent, the detections are transformed to global coordinates, and a map of the environment is continuously built. As the focus of this paper is to learn coverage paths, and not to solve the localization problem, we assume known pose up to a certain noise level. However, our method may be extended to the case with unknown pose with the use of an off-the-shelf SLAM method. Finally, while our method can be further extended to account for moving obstacles and multi-agent coverage, they are beyond the scope of this paper.
Figure 2: (a) Agent-environment interaction: The observation consists of multi-scale maps from (b) and lidar detections, based on which the model predicts continuous control signals for an agent. (b) Illustration of coverage, obstacle, and frontier maps in multiple scales: This example shows $m=4$ scales with a scale factor of $s=2$. All scales are centered at the agent, and discretized into the same pixel resolution, resulting in the multi-scale maps $M_{c}$, $M_{o}$, and $M_{f}$, of size $8\times 8\times 4$ in this example.
III-B CPP as a Markov Decision Process
@@
We formulate the CPP problem as a partially observable Markov decision process (POMDP), where an agent performs actions in an environment such as to maximize a reward. It is defined by the tuple $(\mathcal{S},\mathcal{A},T,R,\Omega,\mathcal{O})$, where $\mathcal{S}$ is the set of possible states, $\mathcal{A}$ is the set of actions, $T:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,\infty)$ is the state-transition function representing the probability density of the next state $s_{t+1}\in\mathcal{S}$ given the current state $s_{t}\in\mathcal{S}$ and the action $a_{t}\in\mathcal{A}$, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\Omega$ is the set of observations, and $\mathcal{O}:\mathcal{S}\times\mathcal{A}\times\Omega\rightarrow[0,\infty)$ is the observation function representing the probability density of the observation $o_{t}\in\mathcal{O}$ given the state $s_{t}\in\mathcal{S}$ and the previous action $a_{t-1}\in\mathcal{A}$. At each discrete time step $t\in\{1,...,N\}$, the environment is in some state $s_{t}$, in which the agent predicts an action $a_{t}\sim\pi(a_{t}|o_{t})$ according to a policy $\pi$, based on the observation $o_{t}$ of the current state. Subsequently, the environment transitions into a new state $s_{t+1}\sim p(s_{t+1}|s_{t},a_{t})$ according to $T$, while emitting a reward $r_{t+1}=R(s_{t+1},a_{t})$ and a new observation $o_{t+1}\sim p(o_{t+1}|s_{t+1},a_{t})$ according to $\mathcal{O}$. The goal for the agent is to maximize the expected future discounted reward $\mathbb{E}(\sum_{k=t}^{N}\gamma^{k-t}r_{k})$, with discount factor $\gamma$.
@@
In our particular problem, the policy is a neural network that predicts continuous control signals, $a_{t}$, for the agent. The state $s_{t}$ includes the full environment geometry with all obstacles, the set of points that have been covered, as well as the pose of the agent. However, the agent only has information about what it has observed, so the observation $o_{t}$ consists of the part of the environment that has been mapped until time step $t$, along with covered points, agent pose, and sensor data. The reinforcement learning loop is depicted in Fig. 2(a). In the setting with a real robot, the first-order Markovian property is not fulfilled, due to the dynamics including inertia and momentum. The usual approach to augment previous states in the current one becomes infeasible if the motion is implicitly represented by the agent-centered environment maps. Thus higher-order effects lead to model errors that need correction in subsequent steps. We go into further detail in Section IV-B.
III-C Observation Space
To efficiently learn coverage paths, the observation space needs to be mapped into a suitable input feature representation for the neural network policy. To this end, we represent the visited points as a coverage map, and the mapped obstacles and boundary of the area as an obstacle map. The maps are discretized into a 2D grid with sufficiently high resolution to accurately represent the environment. To represent large regions in a scalable manner, we propose to use multi-scale maps, which was necessary for large environments. We make this viable through frontier maps that preserve information about the existence of non-covered space, even in coarse scales.
III-C1 Multi-scale Maps
Inspired by multi-layered maps with varying scale and coarseness levels for search-and-rescue [34], we propose to use multi-scale maps for the coverage and obstacles to solve the scalability issue. We extract multiple local neighborhoods with increasing size and decreasing resolution, keeping the grid size fixed. We start with a local square crop $M^{(1)}$ with side length $d_{1}$ for the smallest and finest scale. The multi-scale map representation $M=\{M^{(i)}\}_{i=1}^{m}$ with $m$ scales is constructed by cropping increasingly larger areas based on a fixed scale factor $s$. Concretely, the side length of map $M^{(i)}$ is $d_{i}=sd_{i-1}$. The resolution for the finest scale is chosen sufficiently high such that the desired level of detail is attained in the nearest vicinity of the agent, allowing it to perform precise local navigation. At the same time, the large-scale maps allow the agent to perform long-term planning, where a high resolution is less important. This multi-scale map representation can completely contain an area of size $d\times d$ in $\mathcal{O}(\log d)$ number of scales. The total number of grid cells is $\mathcal{O}(wh\log d)$, where $w$ and $h$ are the fixed width and height of the grids, and do not depend on $d$. This is a significant improvement over a single fixed-resolution map with $\mathcal{O}(d^{2})$ grid cells. For the observation space, we use a multi-scale map $M_{c}$ for the coverage and $M_{o}$ for the obstacles. These are illustrated in Fig. 2(b).
III-C2 Frontier Maps
When the closest vicinity is covered, the agent needs to make a decision where to explore next. However, the information in the low-resolution large-scale maps may be insufficient. For example, consider an obstacle-cluttered region where the obstacle boundaries have been mapped. A low coverage could either mean that some parts are non-covered free space, or that they are part of the interior of an obstacle. These cases cannot be distinguished if the resolution is too low. As a solution to this problem, we propose to encode a multi-scale frontier map $M_{f}$, which we define in the following way. In the finest scale, a non-obstacle grid cell that has not been visited is a frontier point if any of its neighbours have been visited. Thus, a frontier point is non-visited free space that is reachable from the covered area. A region where the entire free space has been visited does not induce any frontier points. In the coarser scales, a grid cell is a frontier cell if and only if it contains at least one frontier point. In this way, the existence of frontier points persists through scales. Thus, regions with non-covered free space can be deduced in any scale, based on this multi-scale frontier map representation.
III-C3 Egocentric Maps
As the movement is made relative to the agent, its pose needs to be related to the map of the environment. Following Chen et al. [41], we use egocentric maps which encode the pose by representing the maps in the coordinate frame of the agent. Each multi-scale map is aligned such that the agent is in the center of the map, facing upwards. This allows the agent to easily map observations to movement actions, instead of having to learn an additional mapping from a separate feature representation of its position, such as a 2D one-hot map [42].
III-C4 Sensor Observations
To react on short-term obstacle detections, we include the sensor data in the input feature representation. The depth measurements from the lidar sensor are normalized to $[0,1]$ based on its maximum range, and concatenated into a vector $S$. Note, however, that this assumes a fixed set if lidar parameters, i.e. for the range, field-of-view, and number of rays.
III-D Action Space
We let the model directly predict the control signals for the agent. This allows it to adapt to specific environment characteristics, while avoiding a constrained path space, different from a discrete action space. We consider an agent that independently controls the linear velocity $v$ and the angular velocity $\omega$, although the action space may seamlessly be changed to specific vehicle models. To keep a fixed output range, the actions are normalized to [1,1][-1,1] based on the maximum linear and angular velocities, where the sign of the velocities controls the direction of travel and rotation.